sandy ran out at 8mph and rode back at 24mph. if the total trip took 8hrs how far did she run and how far did she ride?
Found 2 solutions by lwsshak3, Alan3354:
Answer by lwsshak3(11628)

sandy ran out at 8mph and rode back at 24mph. if the total trip took 8hrs how far did she run and how far did she ride?
***
let x=time Sandy ran
8-x time Sandy rode
distance=speed*travel time
8x=24(8-x)
8x=192-24x
32x=192
x=6
8-x=2
time Sandy ran=6 hrs
time Sandy rode=2hrs

Answer by Alan3354(69443)

sandy ran out at 8mph and rode back at 24mph. if the total trip took 8hrs how far did she run and how far did she ride?
-----------------
Avg speed for the round trip is
2*8*24/(8 + 24) = 384/32
= 12 mi/hr
-------
12 mi/hr * 8 hrs = 96 miles
She ran 48 miles, rode 48 miles